# Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data

### v2

@@ 4.2 Training the Generative Models and Generating New Designs
Given the trained prediction model, we proceed to generate new designs using the three generative model versions and compare their performances. For illustrative purposes, we focus on generating new designs for the graffiti theme, as it is one of the most popular themes in the data. We present the results obtained for the traditional Chinese theme in Appendix B to show the generalizability of our method to different themes.
Baseline model: DCGAN without consumer preferences
@@
We produce output images with dimensions of $3\times 128\times 128$. The latent dimensionality is set to $100$, as is the standard value for the DCGAN. We use a batch size of $32$ for training the model, with the learning rate set to $2\times e^{-4}$, and the sizes of the feature maps in the generator and discriminator are both 64. To obtain stable results, we train the DCGAN for $5000$ epochs, implementing $275,000$ iterations in total.
Enhanced model: CcGAN with internal consumer preferences
@@
To train the CcGAN with internal consumer preferences, we first use Prediction Model 1 to calculate the aggregate popularity levels for each of the 1388 existing templates. One caveat is that the prediction model requires the display rankings of the templates as inputs. We observe the rankings only for the internal templates but not for the external templates, as the external templates are from UGC websites and are not displayed in the system of the firm. We address this issue by simulating rankings for the external templates based on the observed empirical distribution of the rankings of the internal templates. The remainder of the procedure for calculating the popularity levels follows that discussed in Section 2.3.
Training the CcGAN requires a sufficient number of images within each class label. In our case, the popularity level takes continuous values, so the number of images with exactly the same popularity level is small. Therefore, we group the images with similar popularity levels into bins and let the images belonging to the same bin share the same popularity label. We focus on the images with popularity labels within the range of 0.4 to 0.65 since the majority of the images fall within this range. We limit the maximum number of images per label to 200 to produce a more balanced distribution. This results in a dataset containing 1317 images across 11 popularity labels. Figure A.3 in the appendix presents histograms for both the original popularity levels and the binned popularity levels. Finally, to facilitate comparisons across generative models, we normalize each label by dividing it by the maximum label value, resulting in 11 normalized labels ranging from 0.6154 to 1. These normalized labels are then used as continuous labels to guide the new design generation process.
@@
We use the HVDL as the discriminator loss, where the specific loss functions are defined in Equations 2.4 and 2.5. Regarding the hyperparameters of the HVDL, we use the rule-of-thumb formulae to select them, i.e., $\sigma\approx 0.028$ and $\kappa\approx 0.192$. The CcGAN is trained for 20,000 iterations with a constant learning rate of $10^{-4}$ and a batch size of 64. The dimensions of the latent GAN are set to 256, and the discriminator updates twice in one iteration. To solve the small sample size problem, we adopt the DiffAugment method during the training process, similar to what we did in the DCGAN case. After training the CcGAN, we instruct the generator to produce new template designs with different popularity labels, where label=0.6154 indicates the least-preferred class, and label=1 indicates the most-preferred class.
Following a procedure similar to that used for training the enhanced model, we use Prediction Model 2 to calculate popularity levels for the existing templates. We focus on images with popularity labels between 0.65 and 0.9 since the majority of the images fall into this range, and we cap the maximum number of images per label at 200 for balance purposes. This process results in 1283 images, which are then categorized into 10 labels with normalized label values between 0.7429 and 1. Figure A.4 in the Appendix illustrates the original and grouped popularity levels. The hyperparameters of the HVDL, $\sigma\approx 0.019$ and $\kappa\approx 0.143$, follow the rule-of-thumb formulae.
5 Results
In this section, we provide the results of our empirical application involving the unmanned photo gallery. Section 5.1 presents performance tests concerning the prediction models, including the prediction model with internal preferences only and the prediction model with both internal and external preferences. Section 5.2 shows the validity test conducted for the three generative models, including the DCGAN, the CcGAN model with internal preferences, and the CcGAN model with both internal and external preferences. Section 5.3 compares the performances of the three generative models.
5.1 Performance of the Prediction Models
@@
In this section, we present performance tests concerning our prediction models. As the prediction models are trained on templates from all themes, the results are also obtained on all the themes.
Predicting actual consumer choices
We first show that our prediction model is able to predict observed consumer choices. Figure 6 presents examples of the predicted and actual choices of two consumers in the held-out sample. The first column shows the model-predicted consumer choices, which are defined as having predicted selection probabilities exceeding $0.5$. The second column shows the actual choices and the photos taken by the consumer. We can see that the actual choices are among the choices predicted by the model, but some predicted templates were not chosen by the consumers (i.e., false positives). [^2] The false positives in this illustrative test are not a major concern because they change with the threshold that we set. Here, we set a relative conservative threshold of $0.5$ by defining a model-predicted choice as having a predicted selection probability exceeding $0.5$. Setting a larger threshold will reduce the number of false positives. In the generative model, we use the continuous predicted choice probabilities as popularity labels, which are not affected by the threshold we set here. These examples provide direct evidence that our model can capture consumers’ heterogeneous and nonlinear preferences for the template designs fairly well.
Figure 6: Examples of Predicted and Actual Consumer Choices
Consumer Templates with predicted probabilities $>0.5$ Templates chosen by the consumer
Consumer 1 Refer to caption Refer to caption
Consumer 2 Refer to caption Refer to caption
@@
Figure 7 presents the observed and predicted numbers of adoptions for three internal templates in the test set. We can see that the true and predicted numbers of adoptions are quite close to each other, suggesting that our prediction model is able to capture consumer preferences for different template designs.
@@
Design popularity prediction
@@
Prediction accuracy tests against the benchmark models
Next, we conduct prediction accuracy tests for Prediction Model 1 and Prediction Model 2 on the test set.
In the first row of Table 1, Prediction Model 1 has an overall accuracy rate of $0.793$. Given the unbalanced nature of our sample, another important performance measure is the balanced accuracy rate. The balanced accuracy rate of our model is $0.791$, which is almost as high as the overall accuracy rate. We further check the false-negative rate (FNR) and the false-positive rate (FPR) of our model and obtain an FNR of 0.212 and an FPR of $0.206$. These prediction test results show that our prediction model is able to use consumers’ facial traits to predict their template choices at a reasonably high accuracy level.
We also find that incorporating external preferences can help increase the overall prediction accuracy of the model. As shown in the second row of Table 1, the overall accuracy rate of Prediction Model 2 increases to 0.806, and its balanced accuracy rate is even higher, at 0.825. Both the false-negative rate and the false-positive rate are lower.
Finally, we compare our method with the logistic lasso model, which is a commonly used benchmark for binary outcome models. We choose logistic regression with lasso regularization for two main reasons. First, logistic regression is a prevalent linear and parsimonious model for classification tasks. Second, the lasso model is suitable for our context because it helps address the high dimensionality problem of our input variables (i.e., 1001+128+1) by automatically selecting relevant features and penalizing the inclusion of numerous trivial predictors.
Table 1 compares the results of the two models. Our proposed model outperforms the benchmark model with higher accuracy and balanced accuracy rates. This confirms the superiority of our proposed model in terms of predicting consumer template choices.
@@
Note: This table presents the prediction accuracies, balanced accuracy rates, false-negative rates, and false-positive rates achieved on the test set using Prediction Model 1, Prediction Model 2, and the benchmark logistic lasso model. Prediction Model 1 is trained on the internal preferences only. Prediction Model 2 is trained on both the internal and external preferences. The benchmark model is trained on the internal preferences in the first row and both the internal and external preferences in the second row.
5.2 Validity of the Generative Models
After presenting the performance test for the prediction models, we now implement performance tests for the generative models. In this subsection, we demonstrate the face validity of the generative models by showing their ability to generate meaningful product designs with certain popularity levels. In the next subsection, we compare the performances of the generative models by comparing their abilities to generate product designs that are more likely to be preferred by consumers.
@@ 5.2 Validity of the Generative Models
We first show that our generative models are able to generate meaningful new designs. Intuitively, traditional GAN models can use a set of dog images to produce a new set of dog images. In our case, we need to show that our new model-generated designs are derived from the same theme as that of the input designs (i.e., graffiti). We start with a simple visual assessment. Figure 9(b) compares the real images and the new images generated by the trained DCGAN. We can see that the generated images are similar in style to the real images. At the same time, the generated images exhibit sufficient style variations, which guarantees that no modal collapse occurs during the training process.
@@
In addition to this simple visual examination, we use a pretrained image classifier, ResNet18 (He et al.,, 2016), to conduct transfer learning to validate the meaningfulness of our DCGAN-generated images. Intuitively, if the generated images are indeed from the same theme as that of the real images (i.e., graffiti), ResNet18 should be able to classify them into the same category as that of the real images. First, we retrain the classifier to recognize graffiti and nongraffiti images using 100 real graffiti images (both internal and external, with $\text{label}=1$) and 70 real nongraffiti images (internal, with $\text{label}=0$) as inputs. Second, we test the validity of the retrained classifier on a test set consisting of real images. As shown in Table 2, the classifier has an accuracy rate of 97.2% when classifying the images into graffiti and nongraffiti categories. The false-negative rate and false-positive rate are both very low on the test set, further validating the ability of the classifier to predict the graffiti category. Finally, we use this retrained classifier to predict the labels for the set of generated graffiti images and present the hit rate in the last column of Table 2. Here, the hit rate is defined as the percentage of the generated images that are predicted to be in the graffiti category (i.e., $\text{label}=1$). We find that the hit rate is 99%, meaning that almost all the generated images are classified into the graffiti category by the classifier. This means that our generative model is able to produce meaningful designs with high fidelity.
Table 2: Image Classification Test Concerning the Generated Images
Test Set of Real Graffiti Images Generated Images
Classification Model Accuracy FNR FPR Hit Rate
ResNet18 Transfer Learning 0.972 0.024 0.034 0.99
@@
One advantage of our proposed framework is that it can be instructed to generate designs with a certain popularity level. This contrasts with traditional GAN models, which cannot be instructed to create images of a specific type. We start with the CcGAN model that incorporates internal consumer preferences. To demonstrate its ability to generate new designs with different popularity levels, we instruct it to generate new designs from the least-preferred class and the most-preferred class. The first row of Figure 10 presents the generated images. The images derived from both classes have graffiti styles, which means that the model passes the face validity check. Moreover, it is apparent that the generated images differ in style across the two classes, with the least-preferred class being darker in color, and the most-preferred class being brighter. This pattern is consistent with the pattern observed in the internal images.
@@
Similarly, we present the new designs generated for the least-preferred class and the most-preferred class by the CcGAN model with both internal and external preferences. As shown in the second row of Figure 10, the generated images differ in style across the two classes. The difference is more salient than in the case in which only internal preferences are incorporated. The least-preferred class is much darker in color, and the most-preferred class is brighter and has more blue/reddish colors. The image patterns in the most-preferred class also more exemplify the graffiti theme than do those in the least-preferred class. This means that the external preferences help generate images with higher quality in terms of mimicking the graffiti features in the most-preferred class.
As our CcGAN models utilize continuous popularity labels, they are able to generate a range of images with continuous popularity levels. Figure 11 displays the progression of the generated images ranging from the lowest popularity level to the highest popularity level. The top row shows the images generated solely with the internal preferences, while the bottom row shows the images generated with both internal and external preferences. A noticeable trend is the gradual change in the style of the generated images as the popularity level increases in both models. Images with higher popularity levels have brighter colors and clearer patterns, which resemble the observed patterns among the real images.
@@
In summary, the DCGAN, the CcGAN with internal preferences, and the CcGAN with both internal and external preferences all generate meaningful graffiti templates that have similar styles to those of the internal templates. Furthermore, the CcGAN models are able to generate new templates with different popularity levels. This ability is stronger for the CcGAN with both internal and external preferences, highlighting the benefits of incorporating external preferences into the generation task.
5.3 Performance of the Generative Models
After establishing the validity of the generative models in terms of their ability to generate new designs in a controlled manner, we now compare their performances with respect to generating new appealing designs that are likely to be preferred by consumers, which is the main goal of the proposed framework.
Performance based on model-predicted popularity
Given the newly generated designs, we assess how attractive they are to consumers and how likely consumers are to choose them. Our prediction model is trained to predict consumer choices, so we can use our trained prediction model to calculate consumers’ choice probabilities for the new designs generated by the three generative models.
Table 3 shows the average predicted choice probabilities of the internal real templates, DCGAN-generated templates, and CcGAN-generated templates (from the most preferred class). The comparison illustrates whether the generative models can generate new designs that outperform the existing offerings of the company. It also illustrates which generative model can generate the most appealing new designs. As we have two prediction models, we use Prediction Model 1 to evaluate the attractiveness of the templates in the first row and use Prediction Model 2 in the second row. To maintain consistency between the prediction and generation models, we use Prediction Model 1 (Prediction Model 2) to evaluate the performance of the CcGAN with internal preferences only (with internal and external preferences). To facilitate the comparison, we normalize the choice probabilities of the real internal images to 1 in the first column and present the percentage changes relative to the first column in the other columns.
The results show that the new DCGAN-generated designs do not perform well when evaluated by Prediction Model 1. Intuitively, the traditional GAN method (i.e., the DCGAN) only incorporates external templates when generating new templates. Although the external templates help bring new design features to the DCGAN-generated templates, Prediction Model 1 cannot recognize the attractiveness of such new features because it learns only consumer preferences on internal templates. The external features that are different from the popular internal features are treated as unpopular by Prediction Model 1. Therefore, the performance of the DCGAN model is even worse than that attained on the internal real images (-4.43%).
The CcGAN model with internal preferences exhibits better performance by using the internal preferences to guide the generative model, allowing it to generate templates that are more appealing to internal consumers. However, this improvement is limited (0.72%) because Prediction Model 1 is trained only on internal preferences and cannot fully recognize the out-of-the-box features derived from the external templates.
@@ 5.3 Performance of the Generative Models
Further incorporating external preference information allows the CcGAN to fully leverage the appealing features of the external templates and substantially improve the performance of the generative model. The incorporation of external preference information allows Prediction Model 2 to learn the types of features that consumers prefer in the external templates. The underlying assumption is that external consumers with similar appearances to those of the internal consumers have similar preferences, so the preference information of the external consumers can be used to infer the preferences of the internal consumers. Prediction Model 2 is thus able to tell which external features might be appealing to the internal consumers, guiding the CcGAN to incorporate popular features acquired from both internal and external sources and leading to the largest model performance improvement (5.37%). The performance of the DCGAN is also better (3.54%) when evaluated by Prediction Model 2 because this model can now recognize popular features derived from external sources in the DCGAN-generated templates.
@@
Evaluation Methods Internal DCGAN CcGAN CcGAN
Real Images (Internal) (Internal and External)
Prediction Model 1 1.00 -4.43% 0.72% N.A.
Prediction Model 2 1.00 3.54% N.A. 5.37%

### v3

@@ Introduction
Given the trained prediction model, we proceed to training the generator models and generate new designs. While we train the prediction model using samples from all template themes, we train separate generative models for different themes. For illustrative purposes, we present the generative results for the graffiti theme, as it is one of the most popular themes in the data. We present the results for other themes (e.g., the traditional Chinese theme) in the online appendix to show the generalizability of our method to different themes.
@@
We produce output images with dimensions of $128\times 128$. The latent dimensionality is set to $100$, as is the standard value for the DCGAN. We use a batch size of $32$ for training the model, with the learning rate set to $2\times e^{-4}$, and the sizes of the feature maps in the generator and discriminator are both 64. To obtain stable results, we train the DCGAN for $5000$ epochs, implementing $275,000$ iterations in total.
@@
To train the CcGAN with internal consumer preferences, we start by using Prediction Model 1 to calculate aggregate popularity levels for the 1388 existing templates. This model requires display rankings as inputs, which are only available for internal templates because external templates from UGC websites aren’t displayed in the firm’s system. We address this issue by simulating rankings for the external templates based on the empirical distribution of internal template rankings. To compare generative models, we normalize each popularity level by dividing it by the maximum value within each model, resulting in normalized levels from 0 to 1, which serve as continuous input labels for the generator.
@@
We use the HVDL as the discriminator loss, where the specific loss functions are defined in Equations 4 and 5. Regarding the hyperparameters of the HVDL, we use the rule-of-thumb formulae to select them, i.e., $\sigma\approx 0.028$ and $\kappa\approx 0.192$. The label weight parameter $k=10$. The CcGAN is trained for 20,000 iterations with a constant learning rate of $10^{-4}$ and a batch size of 64. The dimensions of the latent GAN are set to 256, and the discriminator updates twice in one iteration. To solve the small sample size problem, we adopt the DiffAugment method during the training process, similar to what we did in the DCGAN case. After training the CcGAN, we instruct the generator to produce new template designs with different popularity labels, where label=1 indicates the most-preferred class.
To train the CcGAN with internal and external preferences, we employ Prediction Model 2 to calculate popularity levels for the existing templates and use the normalized popularity labels as inputs to the generator. The hyperparameters of the HVDL, $\sigma\approx 0.019$ and $\kappa\approx 0.143$, follow the rule-of-thumb formulae. The label weight parameter $k=10$.
Prediction Model Performance
@@
This section presents performance tests for our prediction models, which are trained and evaluated on templates from all themes.
Predicting consumer choices
@@
We first show that our prediction model is able to predict observed consumer choices. Figure 5 presents the observed and predicted numbers of adoptions for three internal templates in the test set. We can see that the true and predicted numbers of adoptions are quite close to each other, suggesting that our prediction model is able to capture consumer preferences for different template designs. Our model is also able to predict choices of individual consumers, which we show in the online appendix.
@@
Predicting design popularity
@@
Prediction accuracy tests against benchmark models
Next, we conduct prediction accuracy tests for Prediction Models 1 and 2 on the test set. As shown in Table 1, Prediction Model 1 achieves an overall accuracy of $0.793$ and a balanced accuracy of $0.791$. The model’s false-negative rate (FNR) the false-positive rate (FPR) are both low, indicating a reasonably high accuracy in predicting consumers’ template choices based on facial traits. Incorporating external preferences further improves the model’s performance. Prediction Model 2’s overall accuracy increases to $0.806$, with a balanced accuracy of $0.825$, and both FNR and FPR are lower. We also compare our method with the logistic lasso model, a common benchmark for handling binary outcomes and high-dimensional (1001+128+1 in our case) input data. Our proposed model outperforms the benchmark, confirming its superiority in predicting consumer template choices.
@@
Note: This table presents the prediction accuracies, balanced accuracy rates, false-negative rates, and false-positive rates for the test set using three models: Prediction Model 1 (trained on internal preferences), Prediction Model 2 (trained on both internal and external preferences), and a benchmark logistic lasso model. The benchmark model is evaluated with internal preferences in the first row and both internal and external preferences in the second row.
Generative Model Validity Checks
@@
We first demonstrate the face validity of the generative models by showing their ability to generate meaningful product designs. Specifically, we need to show that our new model-generated designs are derived from the same theme as that of the input designs (i.e., graffiti). We start with a simple visual assessment. Figure 7(b) compares the real images and the new images generated by the trained DCGAN. We can see that the generated images are similar in style to the real images. At the same time, the generated images exhibit sufficient style variations, which guarantees that no modal collapse occurs during the training process.
@@
In addition to this simple visual examination, we use a pretrained image classifier, ResNet18 (He et al.,, 2016), to conduct transfer learning to validate the meaningfulness of our DCGAN-generated images. Intuitively, if the generated images are indeed from the same theme as that of the real images (i.e., graffiti), ResNet18 should be able to classify them into the same category as that of the real images. First, we retrain the classifier to recognize graffiti and nongraffiti images using 100 real graffiti images (both internal and external, with $\text{label}=1$) and 70 real nongraffiti images (internal, with $\text{label}=0$) as inputs. Second, we test the validity of the retrained classifier on a test set consisting of real images. The classifier has an accuracy rate of 97.2% when classifying the images into graffiti and nongraffiti categories. The false-negative rate and false-positive rate are both very low (2.4% and 3.4%) on the test set, further validating the ability of the classifier to predict the graffiti category. Finally, we use this retrained classifier to predict the labels for the set of generated graffiti images and present the hit rate, defined as the percentage of the generated images that are predicted to be in the graffiti category (i.e., $\text{label}=1$). We find that the hit rate is 99%, meaning that almost all the generated images are classified into the graffiti category by the classifier. This means that our generative model is able to produce meaningful designs with high fidelity.
@@
Our proposed framework allows for generating designs with specific popularity levels, unlike traditional GAN models. We use the CcGAN model with internal consumer preferences to generating new designs from the least-preferred and most-preferred classes. The first row of Figure 8 shows that both classes produce graffiti-style images, passing the face validity check. The least-preferred class features simpler shapes and patterns, while the most-preferred class has more intricate designs, consistent with internal images.
Additionally, we generate new designs using the CcGAN model with both internal and external preferences. The second row of Figure 8 shows more pronounced style differences between the classes. The least-preferred class exhibits simpler colors and shapes, while the most-preferred class displays more intricate and colorful designs, better exemplifying the graffiti theme. This indicates that incorporating external preferences enhances the quality of generated images in the most-preferred class.
@@
As our CcGAN models utilize continuous popularity labels, they are able to generate a range of images with continuous popularity levels. Figure 9 displays the progression of the generated images ranging from the lowest to the highest popularity level. The top row shows the images generated by CcGAN with the internal preferences, while the bottom row shows the images generated by CcGAN with both internal and external preferences. The style of the generated images gradually changes as the popularity level increases. Images with higher popularity levels have more sophisticated and clearer patterns, especially for the images generated by CcGAN with both preferences.
@@
Generative Model Performance Tests
After validating the generative models’ capability to controllably generate designs, we now assess their effectiveness in generating designs favored by consumers, which is the primary objective of our framework. We conduct three types of performance tests: a distance-based metric comparing designs to known popular ones, an online consumer survey, and choice probability assessments using prediction models.
Distance-based metric
We propose a distance-based metric as an indicator of popularity. The core idea is that popular templates share common features, suggesting that consumers tend to favor certain types of templates. By measuring the similarity between newly generated images and existing popular templates, we can determine their potential appeal. If the new templates closely resemble the popular ones, they are more likely to be attractive to consumers. Additionally, if they differ significantly from unpopular templates, their chances of being well-received also increase. The metric provides an independent assessment of model performance because it does not rely on any model within our framework. Instead, it leverages the intrinsic similarity among the image features to suggest popularity.
We start by identifying the most and least popular templates among existing internal and external templates, which serve as the targets for measuring the distances. For internal templates, popularity is determined by the number of consumers selecting each template—the most popular template is chosen by the largest number, while the least popular is chosen by the smallest. For external templates, direct consumer choice data isn’t available since each external image corresponds to a unique consumer’s choice of a unique template. Instead, we assign a popularity score based on visual similarity to frequently observed patterns in the external dataset. Patterns that are repeatedly seen among external templates indicate popularity among external consumers. To identify these popular patterns, we use the Caffe BVLC reference model to obtain feature embeddings for all external templates to capture their semantic and visual characteristics. These embeddings are clustered using KMeans into ten clusters, with the cluster centers representing common or popular visual archetypes. An external template’s popularity score is calculated by computing the cosine similarity between its embedding and each cluster center, averaging these similarities. This score reflects the image’s alignment with popular patterns — images with higher scores are more similar to common archetypes and are thus deemed more popular, while those with lower scores deviate from typical patterns.
The calculation procedure of the distance-based metric is as follows. Denote the most popular internal and external templates $j\in\mathcal{J}$ and the least popular internal and external templates $k\in\mathcal{K}$. First, we use the Caffe BVLC reference model to obtain feature embeddings of the new templates generated by different generative models and the most and least popular existing templates. Second, we calculate the distance between each generated template $i$ and each of the most and least popular existing templates. We further normalize the distances to be within the range of [0, 1]. Let $d_{ij}$ and $d_{ik}$ denote the normalized distances between template $i$ and the most popular template $j$ and the least popular template $k$, respectively. Finally, we calculate the distance-based metric for each generated template and average the results over all templates generated from the same generative model to obtain a measure for each generative model. To summarize, the distance-based metric for each generative model is defined as
where $N$ denotes the total number of templates generated by each generative model, $\mathcal{J}$ denotes the set of the most popular internal and external templates, and $\mathcal{K}$ denotes the set of the least popular internal and external templates. This distance-based metric decreases when the generated templates are more (less) similar to the most (least) popular existing templates. The smaller the distance-based metric is, the better.
We calculate the distance-based metric for four sets of images: 1) the internal original templates; 2) 100 new templates generated by DCGAN; 3) 100 new templates generated by CcGAN with internal preference from the most-preferred class; 4) 100 new templates generated by CcGAN with internal and external preferences from the most-preferred class. Table 2 presents the mean and median distance-based metrics for each set of images.
Table 2: Distance-Based Metric: Mean and Median
Statistics Internal original images DCGAN CcGAN with internal preference CcGAN with internal and external preferences (no label weights) CcGAN with internal and external preferences
Mean 1.32 1.26 1.25 1.18 1.14
Median 1.38 1.25 1.22 1.19 1.13
Note: This table presents the mean and median distance-based metrics for images generated by different GAN models. Smaller values indicate higher popularity.
We find that CcGAN with internal and external preferences performs the best, followed by CcGAN with internal preference and DCGAN. All generative models outperform the original internal templates provided by the company. Specifically, DCGAN (Column 2 of Table 2) performs the worst among the generative models because it does not account for consumer preferences. CcGAN with internal preference (Column 3 of Table 2) performs better as it is guided by Prediction Model 1, which captures internal preferences, enabling it to generate templates more appealing to internal consumers. However, this improvement is limited because Prediction Model 1 is trained solely on internal preferences and fails to fully recognize the appeal of new features from external templates. Consequently, any novel features from external templates that differ from popular internal features are perceived as unpopular by Prediction Model 1.
@@
CcGAN with both internal and external preferences (Column 5 of Table 2) performs the best by integrating external preference information through Prediction Model 2. This integration allows CcGAN to fully exploit appealing features from external templates. Specifically, Prediction Model 2 learns the types of features consumers prefer in external templates. The underlying assumption is that external consumers who resemble internal consumers have similar preferences, so external consumer preferences can inform those of internal consumers. Consequently, Prediction Model 2 can identify which external features might appeal to internal consumers, guiding CcGAN to integrate popular features from both internal and external sources, leading to the greatest improvement in model performance.
It is noteworthy that the label weights in the loss functions (Equations 4 and 5) play a crucial role in enhancing the attractiveness of generated templates. We modify the CcGAN from Ding et al., (2020) by incorporating label weights to account for the ordinal nature of labels in our context. To demonstrate their impact, we trained the CcGAN without label weights. Figure 10 shows sample templates from this version, which appear less sophisticated and more blurred compared to those in Figure 8. The distance-based metric for this CcGAN without label weights (Column 4 of Table 2) indicates worse performance compared to the full model with label weights (Column 5 of Table 2). These findings underscore that incorporating label weights significantly enhances the quality of the generated images.
Figure 10: CcGAN-Generated Images without Label Weights
CcGAN (internal preferences) without label weights
Refer to caption Refer to caption
(a) Least-Preferred Class (b) Most-Preferred Class
CcGAN (internal and external preferences) without label weights
Refer to caption Refer to caption
(a) Least-Preferred Class (b) Most-Preferred Class
Overall, the results reveal three key insights: (1) Integrating consumer preference information enhances performance, with CcGAN consistently outperforming DCGAN. (2) Including external consumer preferences is essential for improving image quality, as it helps the predictor model identify and incorporate appealing external features. (3) Adapting CcGAN to recognize the ordinal nature of labels is crucial for generating high-quality designs.
Online consumer survey
We further obtain direct consumer evaluation on the generated designs through an online survey on Credamo, a leading survey platform in China. [^8] https://www.credamo.com/#/dataMarket Mimicking a self-operated photo booth, respondents choose from photo templates, including original internal templates and those generated by the three generative models. As there are many original and newly generated templates, we randomly select 4 generated templates per generative model (denote as A1-A4 for CcGAN with internal and external preferences, B1-B4 for CcGAN with internal preference, C1-C4 for DCGAN) and 2 templates from the 13 internal ones (denoted as D1 and D2). This fixed set of templates is used for the survey.
In the survey, respondents are introduced to a self-operated photo booth concept where they select photo templates as backgrounds for pictures. They are informed that their choices will help design popular backgrounds, aiding photo studios in optimizing services. Respondents imagine entering a self-service photo booth to take personal photos. They are presented with 8 templates (similar number of templates per theme are presented in the real business setting) and asked to choose the 2 most suitable backgrounds. The 8 templates include: 2 from CcGAN with internal and external preferences (randomly from A1-A4), 2 from CcGAN with internal preference (randomly from B1-B4), 2 from DCGAN (randomly selected from C1-C4), and 2 internal original templates (D1-D2). The display order is randomized, and the set of templates varies per respondent. Each respondent participates in one choice occasion. The survey includes attention checks and basic demographic questions, and responses failing attention checks are excluded. The final sample includes 288 respondents. A sample survey is provided in the online appendix.
Table 3 presents the template-level and model-level choice probabilities from the survey data. Template-level choice probability is calculated as the number of times a template is chosen divided by the number of times it is shown. Model-level choice probability is calculated as the number of times templates from a specific model are chosen divided by the number of times they are shown. We find that CcGAN with internal and external preferences performs the best. The survey results also show that the consumer sample in the survey has similar preferences to the internal consumers, as both groups prefer the internal original image D2 over D1, giving us confidence that the online survey reflects consumer evaluations obtained in the photo booth setting.
Table 3: Consumer Survey: Choice Probabilities
Choice probabilities CcGAN with internal and external preferences CcGAN with internal preference DCGAN Internal original images
Model-level 33.0% 23.3% 16.0% 27.8%
Image-level A1: 24.7% B1: 23.5% C1: 19.2% D1: 17.0%
A2: 45.2% B2: 22.7% C2: 11.4% D2: 38.5%
A3: 36.0% B3: 23.7% C3: 18.4%
A4: 26.1% B4: 23.1% C4: 15.0%
Model-predicted popularity
As an alternative performance test, we use our trained prediction model to evaluate the performance of the generative models. We focus on the same four sets of images as in the distance-based metric test. For each image, we calculate consumers’ choice probabilities using our trained prediction models. To maintain consistency between the prediction and generation models, we use Prediction Model 1 (Prediction Model 2) to evaluate the performance of the CcGAN with internal preferences only (with internal and external preferences). Averaging over the choice probabilities of images within each model, we obtain the model-specific average predicted choice probabilities as shown in Table 4(a). Choice probabilities of original internal images are normalized to 1 in the first column, with percentage changes shown in other columns.
The results are consistent with those in the distance-based metric test. As shown in Table 4(a), CcGAN with internal and external preferences shows the largest performance improvement over the internal original images (9.13%) because Prediction Model 2 allows it to fully incorporate popular features from both internal and external sources. CcGAN with internal preference only shows a limited performance improvement (2%) as Prediction Model 1 is trained solely on internal preferences and cannot fully recognize out-of-the-box features from external templates. DCGAN does not perform well. [^9] DCGAN does not perform well when evaluated by Prediction Model 1. Although DCGAN incorporates new design features from external templates, Prediction Model 1 cannot recognize the attractiveness of these new features and treats them as unpopular. Thus, DCGAN-generated templates perform worse than the original internal images (-1.61%). DCGAN performs better (4.47%) when evaluated by Prediction Model 2, as this model can recognize popular features from external sources in the DCGAN-generated templates.
Incorporating label weights is important; CcGANs without label weights (Table 4(b)) show a smaller improvement in predicted choice probabilities compared to CcGANs with label weights (Table 4(a)).
@@
Evaluation Methods Internal DCGAN CcGAN with CcGAN with Internal
Original Images Internal Preference and External Preference
Prediction Model 1 1.00 -1.61% 2.00% N.A.
Prediction Model 2 1.00 4.47% N.A. 9.13%
Evaluation Methods Internal CcGAN with CcGAN with Internal
Original Images Internal Preference and External Preference
(no label weights) (no label weights)
Prediction Model 1 1.00 0.72% N.A.
Prediction Model 2 1.00 N.A. 5.37%